# One-Shot Label-Only Membership Inference Attacks

### v1

@@ Abstract
We introduce One-Shot Label-Only (OSLO) membership inference attacks (MIAs), which accurately infer a given sample’s membership in a target model’s training set with high precision using just a single query, where the target model only returns the predicted hard label. This is in contrast to state-of-the-art label-only attacks which require $\sim 6000$ queries, yet get attack precisions lower than OSLO’s. OSLO leverages transfer-based black-box adversarial attacks. The core idea is that a member sample exhibits more resistance to adversarial perturbations than a non-member. We compare OSLO against state-of-the-art label-only attacks and demonstrate that, despite requiring only one query, our method significantly outperforms previous attacks in terms of precision and true positive rate (TPR) under the same false positive rates (FPR). For example, compared to previous label-only MIAs, OSLO achieves a TPR that is 7$\times$ to 28$\times$ stronger under a 0.1% FPR on CIFAR10 for a ResNet model. We evaluated multiple defense mechanisms against OSLO.
@@ 1 Introduction
Table 1: Comparison to previous label-only attacks. We report the highest attack precision that these attacks can achieve with a recall greater than 1% on CIFAR-10 using ResNet in the table.
Attack method No knowledge about target model architecture Require no auxiliary data Attack
@@
Transfer Attack [6] 50.9% $|\mathcal{D}_{shadow}|$
Data Augmentation [7] 55.7% $\sim 10$
@@
Figure 1: An illustration of OSLO versus the state-of-the-art boundary attack. boundary attack requires querying the target model thousands of thousands of times, whereas OSLO requires only a single query.
We propose novel One-Shot Label-Only (OSLO) MIAs that can infer a sample’s membership in the target model’s training set with just a single query, even when the target model only returns hard labels for input samples. Table 1 compares OSLO to previous label-only MIAs. OSLO is based on constructing transfer-based adversarial example, with just enough magnitude of adversarial perturbation added to each sample to cause misclassification if the sample is non-member (but not enough if it is a member) based on a set of source and validation models, a technique that was initially proposed in [12]. Prior work has considered using a fixed threshold of adversarial perturbation across all samples to infer membership status, resulting in lower precision [7, 6].
We extensively compared OSLO with state-of-the-art label-only methods [7, 6]. Following recent standards, we evaluated the attacks using metrics such as the true positive rate (TPR) under a low false positive rate (FPR) and performed a precision/recall analysis. Our results show that previous label-only MIAs perfom poorly, exhibiting high false positive rates. In contrast, OSLO achieves high precision in identifying members, outperforming prior work by a significant margin. For example, as shown in Figure 2, OSLO is 7$\times$ to 67$\times$ more powerful than other label-only MIAs in terms of TPR under 0.1% FPR across three datasets using a ResNet [13] model. OSLO achieves over 95% precision on all three datasets on ResNet model with a recall greater than 0.01, while the highest precision achieved by other attacks is only 71.4%, 84.2%, and 67.5% respectively. Furthermore, we conduct an detailed comparison of OSLO with previous work, exploring why previous label-only attacks fail to achieve high precision.
@@ 2.2 Membership inference attacks
Membership inference attacks (MIAs) [27, 28, 29, 30, 31, 32] aim to determine whether a specific data sample is a member of the training dataset of a target model. Existing MIAs often assume that the attacker can access the target model’s confidence scores to infer membership. These score-based attacks exploit the difference in confidence scores for training versus non-training data, as models often exhibit higher confidence for samples they have seen during training [1, 4, 5]. Such attacks can be easily defended by limiting access to only the predicted labels, which reduces the amount of information available to the attacker.
@@ 3.1 Problem definition
$H_{0}:(x,y)\in D\quad vs\quad H_{1}:(x,y)\notin D$ (1)
@@ 3.4 Attack method
Surrogate model training
@@
Algorithm 1 Transferable adversarial example generation
Transferable adversarial example generation
@@
In OSLO, the perturbation budget for each sample is adaptive. Specifically, we utilize the Geometry-Aware (GA) framework proposed in [12], which employs a set of validation models to regulate the amount of perturbation added to each sample. The process of adding perturbations is incremental. During this process, a validation model is used to monitor the transferability of the sample. The perturbation addition process is terminated early—referred to as early stopping—when the perturbation added is sufficient for the sample to deceive the validation model, and the confidence of the correct class on the validation model drops below a specified threshold $\tau$:
@@
Through this method, the process can be terminated when the transferability of each sample reaches a certain level, resulting in tailored perturbations for each sample. This can more effectively distinguish the differences of the sample in the in and out worlds.
@@ 4.1 Experimental setup
We utilized three datasets commonly used in prior works [5]: CIFAR-10 [36], CIFAR-100  [36]and SVHN  [37]. For CIFAR-10 and CIFAR-100, we selected 25,000 samples to train the target model under attack. For SVHN, we randomly selected 2,000 samples to train the target model. For each dataset, we trained the surrogate models used for the attack with a set of data that is the same size as the training set of the target model but disjoint.
@@
Metrics We employ the following metrics to evaluate the effectiveness of the attacks: (i) attack TPR and FPR [5]: Carlini et al. suggest that membership inference attacks should be analyzed using TPR and FPR, especially considering that TPR in a low FPR setting is crucial for assessing the success of MIAs. We report the log ROC curve to show the TPR and FPR of the attack, with a focus on the low FPR regime; (ii) attack precision and recall: Attack precision reflects how reliably an attack can infer membership. High-precision attacks allow an attacker to credibly violate the privacy of a portion of the samples. Specifically, precision is defined as the proportion of true members correctly identified among all positively predicted members by an adversary. Meanwhile, recall quantifies the proportion of true members that the adversary has correctly identified out of all actual members. We discarded average-case metrics such as accuracy and AUC, as they have been shown not to accurately reflect the threat of MIAs [5].
@@ 4.2 Results
Figure 2: ROC curves for various label-only attacks on three different datasets on ResNet18
@@
Figure 3: ROC curves for various label-only attacks on three different datasets on DenseNet121
4.2.1 Evaluating attack TPR under Low FPR
As shown in Figure 2, OSLO outperforms all previous label-only MIAs by a large margin. For example, compared to previous attacks, OSLO achieves a TPR that ranges from 7$\times$ to 67$\times$ higher under 0.1% False Positive Rate (FPR), and from 3.3$\times$ to 31.5$\times$ higher under 1% FPR on ResNet, evaluated across three datasets. On DenseNet, the improvements range from 1.25$\times$ to 33$\times$ under 0.1% FPR, and from 2.8$\times$ to 9.5$\times$ under 1% FPR. It can be observed that previous attacks were unsuccessful in the low FPR regime, whereas OSLO can successfully identify some members even at low FPR. For example, on CIFAR-100 using ResNet, at 0.1% FPR regime, 2 of the 5 previous attacks achieved a TPR of zero, the best of the remaining three managed only a 0.3% TPR, while OSLO reached 6.7% TPR. Under 1% FPR, OSLO could even achieve 18.9% TPR, with other attacks achieving at most 2% TPR. These comparisons show that, although OSLO requires only one query to the target model, it can reliably identify part of the members, while previous attacks, despite needing to perform up to thousands of queries, are almost unable to credibly identify members.
@@ 4.2.2 Precision and recall analysis
Figure 4: Precision-Recall curves for various label-only attacks on ResNet18. Each line represents the trade-off between precision and recall for a different method as the attack parameter is varied.
@@
Figure 5: Precision-Recall curves for various label-only attacks on DenseNet121. Each line represents the trade-off between precision and recall for a different method as the attack parameter is varied.
As shown in Figure 4 and Figure 5, OSLO is the only one that effectively trades recall for precision. For example, by trading recall, OSLO achieves over 90% precision across all three datasets with ResNet as the target model. Specifically, on CIFAR-10, CIFAR-100, and SVHN, with recall rates of 5%, 44.7%, and 9% respectively, OSLO attained precisions of 96.2%, 92.0%, and 90.9%. In contrast, the highest precision achieved by other attacks at similar recall levels was 63.2%, 82.5%, and 65.2% respectively. This demonstrates that OSLO can identify members with high precision, a capability that all previous label-only attacks could not match. Previous methods could not improve precision by adjusting hyper-parameters, due to limitations inherent in their designs. We discuss this further in Section 6.1.
@@ 5.2 Effect of using more than one shot
We further report on combining results from the current threshold with those from previous, higher thresholds. For example, ‘three shots‘ represents feeding the current shot and the previous two shots’ generated images to the target model, identifying samples as members if all three adversarial examples fail to fool the target model. As shown in Figure 7, we find that using more shots does not yield a significant improvement. We attribute this to the observation that even with multiple shots, only the shot with the lowest $\tau$ plays a decisive role.
@@ 6 Discussion
6.1 Why OSLO outperform previous approaches
@@ 6.2 Mitigation
In addition to demonstrating the effectiveness of OSLO, we also evaluated six defensive strategies [39, 40, 41] aimed at mitigating membership privacy leakage.
For each defense mechanism, we trained three target models using different hyper-parameters, and the results are presented in Figure 9. It can be observed that confidence alteration methods such as MemGuard[39] are ineffective against OSLO, as expected. Other methods that mitigate memorization require strong regularization to be effective but result in a decrease in the model’s test accuracy.
7 Conclusion
In this paper, we propose OSLO, One-Shot Label-Only Membership Inference Attack (MIAs). OSLO is based on transfer-based adversarial attacks and identifies members by the different perturbations required when a sample is a member versus a non-member. We empirically demonstrate that OSLO not only operates under label-only settings with a single query but also significantly outperforms all previous label-only attacks by a large margin in terms of true positive rate (TPR) under the same false positive rate (FPR) and attack precision. We also discuss why OSLO outperforms existing methods. Furthermore, we evaluate multiple defenses against OSLO, highlighting its robustness.
Limitations.
Despite achieving impressive results using a single query, OSLO requires training surrogate models, which need an auxiliary dataset from the same distribution as the target model’s training set. However, this is a common assumption in MIAs [1, 5]. Furthermore, this assumption may be relaxed by using synthetic data to train the surrogate model, as demonstrated in previous work [1].
@@ Appendix A Appendix
A.1 Adversarial perturbation magnitude for members and non-members

### v2

@@ Abstract
We introduce One-Shot Label-Only (OSLO) membership inference attacks (MIAs), which accurately infer a given sample’s membership in a target model’s training set with high precision using just a single query, where the target model only returns the predicted hard label. This is in contrast to state-of-the-art label-only attacks which require $\sim 6000$ queries, yet get attack precisions lower than OSLO’s. OSLO leverages transfer-based black-box adversarial attacks. The core idea is that a member sample exhibits more resistance to adversarial perturbations than a non-member. We compare OSLO against state-of-the-art label-only attacks and demonstrate that, despite requiring only one query, our method significantly outperforms previous attacks in terms of precision and true positive rate (TPR) under the same false positive rates (FPR). For example, compared to previous label-only MIAs, OSLO achieves a TPR that is at least 7$\times$ higher under a 1% FPR and at least 22$\times$ higher under a 0.1% FPR on CIFAR100 for a ResNet18 model. We evaluated multiple defense mechanisms against OSLO.
@@ 1 Introduction
Figure 1: An illustration of OSLO versus the state-of-the-art boundary attack. Boundary attack requires querying the target model thousands of times, whereas OSLO requires only a single query.
Table 1: Comparison to previous label-only attacks. We report the highest attack precision that these attacks can achieve with a recall greater than 1% on CIFAR-10 using a ResNet18 model.
Attack method Require no knowledge about target model architecture Require no auxiliary data Attack
@@
Transfer Attack [6] 60.9% $|\mathcal{D}_{shadow}|$
Data Augmentation [7] 69.2% $\sim 10$
@@
We propose novel One-Shot Label-Only (OSLO) MIAs that can infer a sample’s membership in the target model’s training set with just a single query, even when the target model only returns hard labels for input samples. Table 1 compares OSLO to previous label-only MIAs. OSLO is based on constructing transfer-based adversarial example, with just enough magnitude of adversarial perturbation added to each sample to cause misclassification if the sample is non-member (but not enough if it is a member) based on a set of source and validation models, a technique that was initially proposed in [12]. Prior work has considered using a fixed threshold of adversarial perturbation across all samples to infer membership status, resulting in a lower precision [7, 6].
We extensively compared OSLO with state-of-the-art label-only methods [7, 6]. Following recent standards, we evaluated the attacks using metrics such as the true positive rate (TPR) under a low false positive rate (FPR) and performed a precision/recall analysis. Our results show that previous label-only MIAs perfom poorly, exhibiting high false positive rates. In contrast, OSLO achieves high precision in identifying members, outperforming prior work by a significant margin. For example, as shown in Figure 2, OSLO is 5$\times$ to 67$\times$ more powerful than other label-only MIAs in terms of TPR under 0.1% FPR across three datasets using a ResNet18 [13] model. OSLO achieves over 95% precision on all three datasets on ResNet18 model with a recall greater than 0.01, while the highest precision achieved by other attacks is only 71.4%, 84.2%, and 67.9% respectively. Furthermore, we conduct a detailed comparison of OSLO with previous work, exploring why previous label-only attacks fail to achieve high precision.
@@ 2.2 Membership inference attacks
Membership inference attacks (MIAs) [27, 28, 29, 30, 31, 32, 33] aim to determine whether a specific data sample is a member of the training dataset of a target model. Existing MIAs often assume that the attacker can access the target model’s confidence scores to infer membership. These score-based attacks exploit the difference in confidence scores for training versus non-training data, as models often exhibit higher confidence for samples they have seen during training [1, 4, 5]. Such attacks can be easily defended by limiting access to only the predicted labels, which reduces the amount of information available to the attacker.
@@
Existing label-only MIAs typically require many queries to be effective [6, 7] and struggle to achieve high TPRs at low FPRs [11]. A recent work YOQO [34] reduced the query budget. However, by design, it cannot adjust the TPR-FPR trade-off and remains ineffective in the low FPR regime, limiting its practicality.
@@ 3.1 Problem definition
$H_{0}:(x,y)\notin D\quad vs\quad H_{1}:(x,y)\in D$ (1)
@@ 3.4 Attack method
Surrogate model training.
@@
Algorithm 1 Transferable adversarial example generation.
Transferable adversarial example generation.
@@
In OSLO, the perturbation budget for each sample is adaptive. Specifically, we utilize the Geometry-Aware (GA) framework proposed in [12], which employs a set of validation models to regulate the amount of perturbation added to each sample. The process of adding perturbations is incremental. During this process, validation models are used to monitor the transferability of the sample. The perturbation addition process is terminated early—referred to as early stopping—when the perturbation added is sufficient for the sample to deceive the validation model, and the confidence of the correct class on the validation model drops below a specified threshold $\tau$:
@@
where $h_{j}(x;\theta)$ denotes the logits for class $j$ produced by the validation model, and $y$ is the ground truth class. Through this method, the process can be terminated when the transferability of each sample reaches a certain level, resulting in tailored perturbations for each sample. This can more effectively distinguish the differences of the sample in the in and out worlds.
@@ 4.1 Experimental setup
We utilized three datasets commonly used in prior works [5]: CIFAR-10 [38], CIFAR-100  [38]and SVHN  [39]. For CIFAR-10 and CIFAR-100, we selected 25,000 samples to train the target model under attack. For SVHN, we randomly selected 2,000 samples to train the target model. For each dataset, we trained the surrogate models used for the attack with a set of data that is the same size as the training set of the target model but disjoint. Further details on dataset splits can be found in Appendix A.1.1.
@@
Baseline Attacks. We compare OSLO against five state-of-the-art label-only MIAs, including the transfer attack [6], three boundary attacks [7, 6], and the data augmentation attack [7]. Detailed settings are provided in Appendix A.1.4. We exclude YOQO [34] from our evaluation, as it is not compatible with our evaluation metrics (see Appendix A.6).
Metrics. We employ the following metrics to evaluate the effectiveness of the attacks: (i) attack TPR and FPR [5]: Carlini et al. suggest that membership inference attacks should be analyzed using TPR and FPR, especially considering that TPR in a low FPR setting is crucial for assessing the success of MIAs. We report the log ROC curve to show the TPR and FPR of the attack, with a focus on the low FPR regime; (ii) attack precision and recall: Attack precision reflects how reliably an attack can infer membership. High-precision attacks allow an attacker to credibly violate the privacy of a portion of the samples. Specifically, precision is defined as the proportion of true members correctly identified among all positively predicted members by an adversary. Meanwhile, recall quantifies the proportion of true members that the adversary has correctly identified out of all actual members. We discarded average-case metrics such as accuracy and AUC, as they have been shown not to accurately reflect the threat of MIAs [5].
@@ 4.2 Results
4.2.1 Evaluating attack TPR under Low FPR
As shown in Figure 2, OSLO outperforms all previous label-only MIAs by a large margin. For example, compared to previous attacks, OSLO achieves a TPR that ranges from 5$\times$ to 67$\times$ higher under 0.1% FPR, and from 3$\times$ to 16$\times$ higher under 1% FPR on ResNet18, evaluated across three datasets. On DenseNet121, the improvements range from 2$\times$ to 12$\times$ under 1% FPR. It can be observed that previous attacks were unsuccessful in the low FPR regime, whereas OSLO can successfully identify some members even at low FPR. For example, on CIFAR-100 using ResNet18, at a 0.1% FPR, the best of the other label-only MIAs achieves only a 0.3% TPR, while OSLO reaches 6.7%. Under a 1% FPR, OSLO achieves a TPR of 18.9%, with other attacks reaching at most 2.7%. These comparisons show that, although OSLO requires only one query to the target model, it can reliably identify part of the members, while previous attacks, despite needing to perform up to thousands of queries, are almost unable to credibly identify members.
@@ 4.2.1 Evaluating attack TPR under Low FPR
Figure 2: ROC curves for various label-only attacks on three different datasets on ResNet18. Each line represents the TPR of an attack under different FPRs, with an emphasis on the low-FPR regime using a logarithmic scale.
@@
Figure 3: ROC curves for various label-only attacks on three different datasets on DenseNet121. Each line represents the TPR of an attack under different FPRs, with an emphasis on the low-FPR regime using a logarithmic scale.
@@ 4.2.2 Precision and recall analysis
As shown in Figure 4 and Figure 5, OSLO is the only one that effectively trades recall for precision. For example, by trading recall, OSLO achieves over 90% precision across all three datasets with ResNet18 as the target model. Specifically, on CIFAR-10, CIFAR-100, and SVHN, with recall rates of 5%, 44.7%, and 9% respectively, OSLO attained precisions of 96.2%, 92.0%, and 90.9%. In contrast, the highest precision achieved by other attacks at similar recall levels was 67.4%, 82.5%, and 65.2% respectively. This demonstrates that OSLO can identify members with high precision, a capability that all previous label-only attacks could not match. Previous methods could not improve precision by adjusting hyper-parameters, due to limitations inherent in their designs. We discuss this further in Section 6.1.
@@
Figure 4: Precision-Recall curves for various label-only attacks on ResNet18. Each line represents the trade-off between precision and recall for an attack as the attack parameter is varied.
@@
Figure 5: Precision-Recall curves for various label-only attacks on DenseNet121. Each line represents the trade-off between precision and recall for an attack as the attack parameter is varied.
@@ 5.1 Effect of using different adversarial attacks
5.2 Effectiveness of OSLO when target and surrogate models use different training algorithms
We assume that the attacker is unaware of the target model’s architecture, using surrogate models with different structures. We further explore scenarios where the surrogate and target models are trained with different optimization algorithms. In particular, we evaluated OSLO on CIFAR-10 using ResNet18, where the target models were trained with SGD, while the surrogate models used Adam. The results are shown in Table 2.
Table 2: OSLO performance on CIFAR-10 with ResNet18 when target and surrogate models are trained using different algorithms.
Target model Source/validation models Attack precision (Recall > 1%) Attack TPR@1% FPR
SGD, lr=0.01 Adam, lr=0.001 95.9% 10.0%
SGD, lr=0.1 Adam, lr=0.001 95.3% 9.2%
Notably, OSLO achieves over 95% precision, with a TPR exceeding 9% at 1% FPR in both settings. These results confirm that OSLO remains effective even when the target models are trained with algorithms different from those of the surrogate models.
5.3 Impact of validation models in OSLO
Table 3: Attack TPR and FPR of OSLO without validation models on CIFAR-10 using ResNet18.
Perturbation budget $\epsilon$ TPR FPR
8/255 20.1% 13%
16/255 2.3% 1.4%
32/255 0.1% 0.1%
OSLO uses validation models to adjust the scale of perturbations added to adversarial examples. To understand the impact of validation models, we conducted an ablation study on CIFAR-10 using ResNet18 without the validation models, where a uniform perturbation budget was applied to all samples. The results are shown in Table 5.3.
These results indicate that without the validation model, applying a uniform perturbation budget across all samples leads to significantly reduced attack effectiveness. For example, at $\epsilon=32/255$, nearly all members and non-members were misclassified, resulting in both TPR and FPR being very low. This highlights the critical role of validation models in effectively calibrating perturbations.
@@ 5.4 Effect of using more than one shot
We further report on combining results from the current threshold with those from previous, higher thresholds. For example, ’three shots’ represents feeding the current shot and the previous two shots’ generated images to the target model, identifying samples as members if all three adversarial examples fail to fool the target model. As shown in Figure 7, we find that using more shots does not yield a significant improvement. We attribute this to the observation that even with multiple shots, only the shot with the lowest $\tau$ plays a decisive role.
@@ 6 Discussion
6.1 Why OSLO outperforms previous approaches
@@ 6.2 Mitigation
In addition to demonstrating the effectiveness of OSLO, we also evaluated six defensive strategies [41, 42, 43] aimed at mitigating membership privacy leakage. These experiments were conducted on the CIFAR-10 dataset using ResNet18 as the target model.
We trained three target models with different hyperparameters for each defense mechanism that requires retraining. Detailed configuration settings are provided in Appendix A.1.5. The results are presented in Figure 9. It can be observed that confidence alteration methods, such as MemGuard [41], are ineffective against OSLO, as expected. Other methods that mitigate memorization require strong regularization to be effective but result in a decrease in the model’s test accuracy. Additionally, we tested OSLO against target models trained with adversarial training (details in Appendix A.5), and observed that while adversarial training reduces OSLO’s effectiveness to some extent, it is not sufficient as a standalone defense.
7 Conclusions and limitations
In this paper, we propose OSLO, One-Shot Label-Only Membership Inference Attack (MIAs). OSLO is based on transfer-based adversarial attacks and identifies members by the different perturbations required when a sample is a member versus a non-member. We empirically demonstrate that OSLO not only operates under label-only settings with a single query but also significantly outperforms all previous label-only attacks by a large margin in terms of true positive rate (TPR) under the same false positive rate (FPR) and attack precision. Thus, OSLO sets a new benchmark for label-only MIAs and can serves as an effective measure of privacy leakage in these settings. We also delve into why OSLO outperforms existing methods. Furthermore, we evaluate multiple defenses against OSLO, highlighting its robustness.
OSLO requires training surrogate models that need an auxiliary dataset from the same distribution as the target model’s training set. Although this is a common assumption in MIAs [1, 5] and can be alleviated by using synthetic data as demonstrated in prior work [1], it remains a limitation.
Acknowledgments
The work was supported in part by the NSF grant 2131910, and by DARPA under Grant DARPA-RA-21-03-09-YFA9-FP-003.
@@ Appendix A Appendix
A.1 Additional Experimental Details
A.1.1 Dataset Split
For CIFAR-10 and CIFAR-100, we used half of the training set (25,000 samples) to train the target model and the remaining half (25,000 samples) to train the source and validation models. The source and validation models were trained on the same data but remained disjoint from the data used to train the target model. Similarly, for SVHN, we created two disjoint subsets of 5,000 samples each, one for training the target model and the other for the surrogate models.
For the evaluation of the attacks, we randomly selected 1,000 samples from both the target model’s training set and the remaining unused samples as target samples. The dataset split is summarized in Table 4.
Table 4: Summary of dataset splits for training and evaluation.
Dataset Target model Source/validation model Target
training training samples
CIFAR-10 25,000 25,000 1,000 (members) + 1,000 (non-members)
CIFAR-100 25,000 25,000 1,000 (members) + 1,000 (non-members)
SVHN 5,000 5,000 1,000 (members) + 1,000 (non-members)
A.1.2 Training setup
A.1.3 OSLO Configuration.
The OSLO setup primarily includes the configuration of source and validation models, as well as the parameters for generating adversarial examples. By default, we use Inception [44], ResNeXt50 [46], and PreActResNet18 [48] as the source model architectures. For validation models, we use DenseNet121 [40], VGG13 [45], and ShuffleNetV2 [47] when ResNet18 [13] is the target model, and ResNet18, VGG13, and ShuffleNetV2 when DenseNet is the target model. We follow the assumption that the attacker has no knowledge of the target model architecture and does not include the same architecture in the source/validation models. Details on the default number of source and validation models are summarized in Table 5.
Table 5: Source and validation model configurations per dataset in the default OSLO setup.
Dataset Source models Validation models Total surrogate models
CIFAR-10 Inception $\times$ 1 DenseNet121 $\times$ 3 12
ResNeXt50 $\times$ 1 VGG13 $\times$ 3
PreActResNet18 $\times$ 1 ShuffleNetV2 $\times$ 3
CIFAR-100 Inception $\times$ 1 DenseNet121 $\times$ 5 18
ResNeXt50 $\times$ 1 VGG13 $\times$ 5
PreActResNet18 $\times$ 1 ShuffleNetV2 $\times$ 5
SVHN Inception $\times$ 1 DenseNet121 $\times$ 1 6
ResNeXt50 $\times$ 1 VGG13 $\times$ 1
PreActResNet18 $\times$ 1 ShuffleNetV2 $\times$ 1
More combinations of source and validation models are discussed in Appendix A.3. For generating adversarial examples, we set the maximum perturbation to $80/255$ and the number of sub-procedures to 80, increasing $\epsilon$ by $1/255$ each iteration.
A.1.4 Baseline attack configuration
We summarize the configurations of the baseline attacks here. For the transfer attack, we use a relabeled dataset of the same size as the target model’s training set and follow the original pipeline [6] to label the shadow dataset and train the shadow model. For boundary attacks based on adversarial perturbation (HopSkipJump and QEBA), we set a query budget of 6,000 per sample. For the Gaussian noise-based boundary attack, we set a query budget of 700. We use rotation and translation to construct augmented samples for the data augmentation attack, following the original method described in the paper [7].
A.1.5 Defense Setup
This section outlines the hyperparameters used for the defenses evaluated against OSLO in Section 6.2. For each defense, three different hyperparameters were used to train the defended models, as detailed in Table 6.
Table 6: Hyperparameter configurations for defenses evaluated against OSLO.
Defense Hyperparameter Values
L2 Regularization Regularization parameter {0.01, 0.005, 0.001}
L1 Regularization Regularization parameter {5e-6, 1e-5, 5e-5}
Adversarial Regularization Alpha {5, 6, 7}
Dropout Dropout rate {0.3, 0.5, 0.7}
DPSGD Norm clipping bound 1.2
Noise multiplier {0.005, 0.01, 0.05}
A.2 Adversarial perturbation magnitude for members and non-members